AT the local smoothie stand, a survey showed that 2/3 of all people preferred strawberry smoothies over the mango smoothies. If 2.400 people responded to the survey, how may prefer the mango smoothies?
Answer #1

Total number of people responded to the survey = 2,400

Number of people preferred strawberry smoothie = 2,400*2/3 = 1,600

Number of people preferred mango smoothie = 2,400 - 1,600 = 800

Hence, 800 people preferred mango smoothies.